OpenStudy (anonymous):

Gloria is digging for fossils at a geological site. She has 120 meters of rope and 4 stakes to mark off a rectangular area. Which set of dimensions will create a rectangle using all the rope Gloria has with her? 12 m × 10 m 50 m × 10 m 60 m × 60 m 45 m × 10 m

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (studygurl14):

Test each one

OpenStudy (skullpatrol):

recall the perimeter of a rectangle formula

OpenStudy (studygurl14):

For example: The first one - 12m x 10 m Perimeter: 2(12) + 2(10) = 44 Too little rope

OpenStudy (skullpatrol):

$$\Huge P = 2l+2w$$

OpenStudy (studygurl14):

So..A cannot be an option. See what I did @beepboop ?

OpenStudy (anonymous):

yes o:

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

So, D

OpenStudy (studygurl14):

not quite. 2(45) + 2(10) = 110

OpenStudy (anonymous):

right xD sorry I'm half-arsing

OpenStudy (studygurl14):

Look at b. 50 x 2 = 100 10 x 2 = 20 100+20=120

OpenStudy (studygurl14):

half-arsing...lol

OpenStudy (anonymous):

B. right

OpenStudy (anonymous):

shhh I got it before I read that ;-;

OpenStudy (anonymous):

I don't wanna get banned for foul language haha I'm having a hard time restraining my fingers

OpenStudy (studygurl14):

lol, yea

OpenStudy (anonymous):

and you can't draw a quadrilateral with no parallel lines and at least one right angle, right? haha

OpenStudy (anonymous):

pretty sure you need parallel lines .-.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

not necessarily....

OpenStudy (studygurl14):

a qudralateral is ANY figure that has four sides

OpenStudy (anonymous):

I can't really think of any that don't have parallel lines xD but thank you!

OpenStudy (studygurl14):

thsese are all quads...

OpenStudy (anonymous):

well that makes sense xD I guess my imagination doesn't work well

OpenStudy (studygurl14):

:)

OpenStudy (studygurl14):

A parallelogram, on the other hand, is a quadrilateral that MUST have two pairs of parallel sides

OpenStudy (studygurl14):

examples

OpenStudy (studygurl14):